OpenStudy (anonymous):

Let and f(x) = x^2 + 6 and g(x) =x+8/x Find ( g o f)(­ -7) A. 295/49 B. 63/55 C. 384/7 D. -57/7

OpenStudy (xapproachesinfinity):

\[gof(-7)=g(f(-7))=g((-7)^2+6)=g(55)=.....\]

OpenStudy (xapproachesinfinity):

finish the work

OpenStudy (xapproachesinfinity):

plug in 55 into g

OpenStudy (anonymous):

When I factored it I got f(−7)=55g

OpenStudy (anonymous):

g o f(−7)=55g @xapproachesinfinity